### v2

# How Semilocal Are Semilocal Density Functional Approximations? —Tackling Self-Interaction Error in One-Electron Systems

Abstract
Self-interaction error (SIE), arising from the imperfect cancellation of the spurious classical Coulomb interaction between an electron and itself, is a persistent challenge in modern density functional approximations. This issue is illustrated using the prototypical one-electron system $H_{2}^{+}$. While significant efforts have been made to eliminate SIE through the development of computationally expensive nonlocal density functionals, it is equally important to explore whether SIE can be mitigated within the framework of more efficient semilocal density functionals. In this study, we present a non-empirical meta-generalized gradient approximation (meta-GGA) that incorporates the Laplacian of the electron density. Our results demonstrate that the meta-GGA significantly reduces SIE, yielding a binding energy curve for $H_{2}^{+}$ that matches the exact solution at equilibrium and improves across a broad range of bond lengths over those of the Perdew-Burke-Ernzerhof (PBE) and strongly-constrained and appropriately-normed (SCAN) semilocal density functionals. This advancement paves the way for further development within the realm of semilocal approximations.
Kohn-Sham density functional theory (KS-DFT) Kohn and Sham (1965) is a widely used quantum mechanical method in condensed-matter physics and quantum chemistry due to its balance of computational efficiency and accuracy. DFT reduces the complexity of many-electron systems by describing their ground-state properties in terms of the electron density, rather than the many-body wavefunction Hohenberg and Kohn (1964), significantly lowering the computational cost. It is exact in theory for a system of electrons under an external potential typically generated by nuclei with the total electronic energy being given by,
$E[n]=T_{s}[n]+E_{H}[n]+E_{en}[n]+E_{xc}[n]$ (1)
where $n(r)$ is the electron density, $T_{s}[n]$ the kinetic energy, $E_{H}[n]$ the Hartree energy, $E_{en}[n]$ the electron-nuclei interaction, and $E_{xc}[n]$ the exchange-correlation energy. However in practice the exchange-correlation (XC) energy functional, $E_{xc}[n]$ = $E_{x}[n]+E_{c}[n]$ has to be approximated as its exact form in terms of $n(r)$ is not explicitly known.
One of the most significant challenges in widely used density functional approximations (DFAs) is the self-interaction error (SIE) Perdew and Zunger (1981), which arises from the inability of these approximations to fully cancel the spurious classical Coulomb interaction of electron with itself. For example, for any one electron system, there is no electron-electron interaction, and thus the correlation energy should be zero, $E_{c}[n]=0$, and the exchange energy should exactly cancel out the Hartree energy, $E_{x}[n]$ = -$E_{H}[n]$. The Hartree-Fock (HF) method satisfies these conditions exactly and proves to be an exact method for describing all one-electron systems.
However, this is typically not the case for DFAs, in particular for the widely used efficient semilocal approximations. SIE in DFAs can lead to inaccuracies in predictions of properties such as bond dissociation energies, transition states, magnetic properties, and band gaps Cohen et al. (2008); Ruzsinszky et al. (2007); Johnson et al. (1994); Hrouda et al. (1997); Patchkovskii et al. (2001); Mori-Sánchez et al. (2008); Toher and Sanvito (2008); Zhang et al. (2020, 2019, 2022); Fitzhugh et al. (2023). Over the years, various methods like Perdew-Zunger self-interaction correction (PZ-SIC) and its variant Perdew and Zunger (1981); Pederson et al. (2014a); Zope et al. (2019); Pederson et al. (2014b), localized orbital scaling correction (LOSC) Li et al. (2018); Mahler et al. (2022), local hybrid density functionals Perdew et al. (2008); Kirkpatrick et al. (2021); Haasler et al. (2020), and other nonlocal density functionals Becke (2013); Janesko (2010); Kong and Proynov (2015) have been developed to reduce or eliminate SIE, offering improvements but often at the cost of increased computational demands due to their nonlocal nature.
Therefore, it is equally important to evaluate how well semilocal density functional approximations (DFAs) handle SIE due to their computational efficiency. This is especially relevant given the success of the Strongly Constrained and Appropriately Normed (SCAN) Sun et al. (2015) functional, which has shown significant improvements over the Perdew-Burke-Ernzerhof (PBE) Perdew et al. (1996) functional across a wide range of materials and properties Sun et al. (2016). SCAN achieves these advancements partially by reducing SIE within the semilocal framework Zhang et al. (2020, 2019, 2022); Fitzhugh et al. (2023), as shown in Figure 1 of the prototypical $H_{2}^{+}$ binding energy curve, which is commonly used to define and illustrate SIE. This raises an intriguing question: can SIE be further minimized while retaining the computational efficiency of semilocal DFAs? Addressing this challenge could lead to the development of even more accurate and efficient functionals, expanding their utility across various applications.
Figure 1: $H_{2}^{+}$ binding energy curves of HF and exchange-only density functionals evaluated on HF orbitals.
In this letter, we present a positive answer to this critical question by introducing a non-empirical semilocal density functional that depends on the Laplacian of the electron density. As shown in Figure 1, this newly designed functional, referred to as RS, significantly reduces SIE, demonstrating clear improvements over other semilocal DFAs like SCAN and PBE in the $H_{2}^{+}$ binding energy curve.
Semilocal DFAs can be written as,
$E_{xc}[n]=\int d^{3}\textbf{r}n\epsilon_{xc}(n,\nabla n,\nabla^{2}n,\tau).$ (2)
In local density approximation (LDA) Von Barth and Hedin (1972); Perdew and Wang (1992), the electron density, $n(\textbf{r})$ is the only ingredient. Generalized gradient approximations (GGA) Perdew et al. (1996); Becke (1988); Perdew et al. (1992) add the electron density gradient, $\nabla n$. Meta-GGAs Perdew et al. (1999); Tao et al. (2003); Sun et al. (2012, 2015); Furness et al. (2020); Mejia-Rodriguez and Trickey (2017); Mejía-Rodríguez and Trickey (2020); Kaplan and Perdew (2022); Lebeda et al. (2024) further include the Laplacian of electron density, $\nabla^{2}n$, and/or the kinetic energy density, $\tau=\frac{1}{2}\sum_{i=1}^{N_{\mathrm{occ}}}|\nabla\phi_{i}|^{2}$, where $\phi_{i}$ are Kohn-Sham orbitals. With more ingredients, semilocal DFAs can satisfy more exact constraints, which are the known properties of the exact exchange-correlation energy Sun et al. (2015); Kaplan et al. (2023).
SCAN is a metaGGA that doesn’t depend on $\nabla^{2}n$ and uses $\tau$ to identify different chemical environments through the variable $\alpha=(\tau-\tau^{W})/\tau^{unif}$ with $\tau^{W}=|\nabla n|^{2}/8n$ and $\tau^{unif}=(3/10)(3\pi^{2})^{2/3}n^{5/3}$ Sun et al. (2013). The $\tau$-dependence also allows SCAN to satisfy all 17 exact constraints suitable for semilocal functionals, including those for iso-orbital (one- and two-electron) systems that are not amenable to GGAs. Moreover, it provides the flexibility to guide the functional between exact constraints using appropriate norms, for example, the hydrogen atom, whose exchange-correlation holes are localized and deep Sun et al. (2015).
Meta-GGAs that depend on $\nabla^{2}n$ can be derived by de-orbitalizing $\tau$-dependent meta-GGAs, by replacing $\tau$ with a function of $\nabla^{2}n$ that offers faster computations. The SCAN-L functional Mejia-Rodriguez and Trickey (2017) is a de-orbitalized version of SCAN, and it demonstrates surprisingly comparable accuracy for many systems and properties Mejia-Rodriguez and Trickey (2017, 2018), including the binding energy curve of $H_{2}^{+}$, as shown in Figure 1. However, since $\alpha$ is only approximated by the de-orbitalized $\alpha_{L}(\nabla^{2}n)$, SCAN-L only approximately satisfies many of the exact constraints that SCAN obeys. For instance, SCAN-L is not self-correlation free in one-electron systems. Furthermore, de-orbitalization does not guarantee that SCAN-L will be accurate for all appropriate norms. As shown in Table 1, SCAN-L underestimates the exchange energy of the hydrogen atom in magnitude.
For one-electron systems, $\tau=\tau^{W}=\frac{|\nabla n|^{2}}{8n}$, and the SCAN meta-GGA loses its dependence on $\tau$ and thus reduces to a GGA. In contrast, SCAN-L remains a meta-GGA, still dependent on $\nabla^{2}n$. Interestingly, the inclusion of this additional ingredient does not improve SCAN-L over SCAN for the binding energy curve of $H_{2}^{+}$, and even worsens its performance for the exchange energies of two analytically known one-electron densities in Table 1, namely H and the Gaussian electron density (G). This raises a question: can SIE be further reduced in one-electron systems by properly incorporating $\nabla^{2}n$ into the functional design, while still satisfying exact constraints and being guided by appropriate norms? Addressing this question would also provide insight into reducing SIE within the semilocal approximation framework, as the $\tau$-dependence could be layered on top of the $\nabla^{2}n$-dependence at the meta-GGA level.
To address this question, we start with the $\nabla^{2}n$-dependent meta-GGA exchange energy of a spin-unpolarized density,
$E_{x}^{MGGA}[n]=\int d^{3}\textbf{r}n\epsilon_{x}^{unif}F_{x}(s,q)$ (3)
where $\epsilon_{x}^{unif}=-(3/4\pi)(3\pi^{2}n)$ is the exchange energy per particle of a uniform electron gas. $F_{x}(s,q)$ is the exchange enhancement factor depending on two dimensionless ingredients, the reduced density gradient $s=|\nabla n|/[2(3\pi^{2})^{1/3}n^{4/3}]$ and the reduced electron density Laplacian $q=\nabla^{2}n/[4(3\pi^{2})^{2/3}n^{5/3}]$. The exchange energy for any spin-polarized density can be found from the exchange energy of a spin-unpolarized density using the exact spin-scaling relationship Oliver and Perdew (1979). We don’t consider the correlation energy functional in this study as $\tau$-dependent metaGGAs can be made one-electron self-correlation free Tao et al. (2003); Sun et al. (2015).
For one-electron systems, the SCAN meta-GGA reduces to a GGA, with
$F_{x}^{SCAN}(s)=1.174\left(1-e^{-\frac{a}{\sqrt{s}}}\right).$ (4)
By design, $F_{x}^{SCAN}\geq 0$, satisfying the exact constraint of exchange energy being negative Perdew and Kurth (2003), $F_{x}^{SCAN}\leq 1.174$, satisfying the tight Lieb-Oxford bound for any one-electron densities Perdew et al. (2014), and $F_{x}^{SCAN}\rightarrow s^{-1/2}$ as $s\rightarrow\infty$, satisfying the finite exchange energy per electron under the one-dimensional, nonuniform coordinate scaling Levy (1991). The parameter $a$ = 4.9479 is then determined by the exchange energy of H. The resulting $F_{x}^{SCAN}(s)$ decreases monotonically with $s$, providing greater enhancement for electron densities with small $s$, which are typically more compact and slowly varying, than those with large $s$ that are dilute and rapidly varying. Such a simple construction works surprisingly well for one-electron systems, improving over PBE, as demonstrated in Table 1 for H and Gaussian electron density and Figure 1 for $H_{2}^{+}$. It is thus desirable to keep this simplicity when introducing the q dependence into $F_{x}(s,q)$.
Table 1: Comparison of exchange energies (in Ha) of different functionals for H and Gaussian electron densities.
System LSDA PBE SCAN SCAN-L RS Exact
Hydrogen -0.2680 -0.3059 -0.3125 -0.3110 -0.3125 -0.3125
Gaussian -0.3410 -0.3819 -0.3975 -0.3965 -0.3989 -0.3989
We therefore design our RS metaGGA by coupling the enhancement factor of SCAN with a function $g(s,q)$,
$F_{x}^{RS}(s,q)=1.174\left(1-e^{-\frac{a}{\sqrt{s}}}\right)\cdot g(s,q),$ (5)
where
$g(s,q)=\left(\frac{1}{1+\ln\left(1+e^{b\cdot(q-q_{0}(s))}\right)}\right).$ (6)
Here, $a$, $b$, and $q_{0}(s)$ are two free parameters and a function of $s$, respectively, which will be determined by appropriate norms to be discussed later. By construction, $0\leq g(s,q)\leq 1$ for any $s\geq 0$ and $-\infty\leq q\leq\infty$. This guarantees that RS satisfies the exact constraints of exchange energy being negative and the tight Lieb-Oxford bound for any one-electron densities.
$g(s,q)$ is also designed to monotonically decrease with $q$, in view of the success of $F_{x}^{SCAN}(s)$ being monotonically decreasing with $s$ which was discussed earlier. This requires $b>0$. According to Bader’s theory of atoms in molecules (AIM) Bader (1985), the behavior of $\nabla^{2}n$ provides critical insights into electronic structure. In regions of high electron density, such as at bond centers for compressed or equilibrium bonds and at nuclei, $\nabla^{2}n$ is negative. Conversely, in regions of low electron density, including inter-shell regions and bond centers of highly stretched bonds, $\nabla^{2}n$ becomes positive. For example, as the bond of $H_{2}^{+}$ is stretched, q at the bond center becomes more positive (see section II in Supplementary Material).
For an electron density under the one-dimensional, nonuniform coordinate scaling, $q\to-\infty$ for some region as the coordinate scaling strength goes to $\infty$, and $q\to\infty$ for the other region (see section III in Supplementary Material). In the region where $q\to-\infty$, $g(s,q)\to 1$, while for the other region, $q\to\infty$, $g(s,q)\to 0$. Therefore, $g(s,q)$ guarantees that $F_{x}^{RS}$ reduces to $F_{x}^{SCAN}$ for the $q\to-\infty$ region, and 0 for the other region, yielding a finite exchange energy for one-electron systems under the one-dimensional, nonuniform coordinate scaling.
The appropriate norms used here to determine $a$, $b$, and $q_{0}(s)$ are H and Gaussian electron density, which have analytic expressions (see section IV of Supplementary Material). Figure 2 (a) shows the system- and spherically-averaged exchange hole Hou et al. (2024a, b) against the inter-electron distance, $u$, for H and Gaussian electron density along with $H_{2}^{+}$ at various bond lengths. In comparison to H, Gaussian electron density as well as $H_{2}^{+}$ with the bond length, $R\leq 1.058$Å have deeper exchange holes, while $H_{2}^{+}$ with the bond length, $R\geq 1.270$Å have shallower ones. It suggests that Gaussian electron density is as qualified for being an appropriate norm as H for semilocal density functionals since Gaussian electron density is also analytically known and not related to bonding. From the comparison between the PBE exchange holes and the exact ones given in section V of Supplementary Material, it is obvious that PBE’s exchange holes are more short-ranged than the exact ones, and become too deep for large bond lengths R of $H_{2}^{+}$. It is also interesting to note that although the exact exchange holes of $H_{2}^{+}$ at equilibrium ($R=1.058$) and Gaussian electron density are deeper than those of H, SCAN predicts an overly negative exchange energy for the former (see Figure 1) and a less negative one for the latter (see Table 1). This suggests that the tailored GGA (eq. 4) may be limited by its variables and could benefit from incorporating $q$ for improved accuracy.
Figure 2: System-and spherically-averaged exact exchange holes plotted against the inter-electron distance $u$ for H, Gaussian electron density, and $H_{2}^{+}$ with various bond lengths.
Figure 3: (a) Contour Plot of the RS enhancement factor against s and q. (b) Enhancement factor of RS as a function of ($q-q_{0}(s)$) for various values of s.
To match the exact exchange energies of the appropriate norms, i.e., H and Gaussian electron density, incorporating q-dependence through $g(s,q)$ in the RS meta-GGA requires careful design of the function $q_{0}(s)$. Because of the spherical symmetry and monotonicity in electron density, q(r) of H and Gaussian electron density can be expressed as functions of s(r), and we denote them as $q_{H}(s)$ and $q_{G}(s)$, plotted in Figure 3 (a) as black and green dashed lines, respectively. For most of the regions, $q_{G}(s)<q_{H}(s)$. $q_{H}(s)$ has an analytical form (see section VI of the Supplementary Material for the derivation),
$q_{H}(s)=s^{2}\left[1-\frac{2}{3\ln{\left((6\pi)^{1/3}s\right)}}\right]$ (7)
Our choice of $q_{0}(s)$ is a modification of $q_{H}(s)$,
$q_{0}(s)=s^{2}\left[1-\frac{2}{3\ln{\left((6\pi)^{1/3}(1+s^{2})^{1/2}\right)}}\right]$ (8)
by turning $s$ in the denominator into $\sqrt{1+s^{2}}$. $q_{0}(s)$ eliminates the singularity of $q_{H}(s)$ and approaches $q_{H}(s)$ asymptotically from above, as shown in Figure 3 (a). This choice of $q_{0}(s)$ provides the flexibility for $g(s,q)\to 1$, thereby restoring $F_{x}^{SCAN}(s)$ for H and Gaussian electron density where $q_{H/G}(s)-q_{0}(s)<0$. For $H_{2}^{+}$, as the bond length $R$ increases, $q$ generally becomes more positive (see section II of the Supplementary Material), and thus $F_{x}^{RS}(s,q)$ tends to decrease as $g(s,q)$ is inversely proportional to q, which is desirable for a less negative exchange energy.
With the designed $q_{0}(s)$, we determine the parameters $a=5.93$ and $b=36.29$ by fitting to the exact exchange energies of H and Gaussian electron density. This parameter set ensures exact matches to the exchange energies of both H and Gaussian electron density, as shown in Table 1. Figure 3(b) illustrates that the enhancement factor decays rapidly with $q$ for a fixed $s$ when crossing $q_{0}(s)$, due to the large value of $b$. As $q_{G}(s)-q_{0}(s)<-0.25$ across most of the $s$-range in Figure 3(a), Gaussian electron density primarily experiences the form of $F_{x}^{SCAN}(s)$. Consequently, the parameter $a=5.93$ is larger than SCAN’s 4.9479 to match the exact exchange energy of Gaussian electron density, leading to an enhancement factor that decays more slowly with $s$ than in SCAN.
Figure 4: (a) Contour Plot of the difference in the RS and SCAN enhancement factors against s and q, overlaid by (s, q) points existing in $H_{2}^{+}$ at the equilibrium bond length ($R=1.058$Å). The inset zooms in the range of 1.4 < s < 2.2. (b) Same as (a) but for $H_{2}^{+}$ with $R=4.76$Å.
Remarkably, this parameter set predicts the total energy of $H_{2}^{+}$ at equilibrium (R = 1.058   Å) to be -0.6026 Hartrees, exactly matching the Hartree-Fock (HF) value. Additionally, the RS binding energy curve aligns perfectly with the HF curve up to slightly beyond $R=1.058\,\text{\AA}$. This consistency likely reflects the fact that the exchange holes of Gaussian electron densities and $H_{2}^{+}$ (for $R\leq 1.058\,\text{\AA}$) are deeper than those of hydrogen. For stretched bond lengths between $R=1\,\text{\AA}$ and $R=3\,\text{\AA}$, the RS binding energy curve shows significant improvement over SCAN and other density functionals. This improvement highlights the importance of incorporating the reduced density Laplacian $q$ in mitigating self-interaction errors.
To understand how RS improves over SCAN for the binding energy curve of $H_{2}^{+}$, Figure  4 plots (s, q) points existing in $H_{2}^{+}$ at equilibrium (R=1.058 Å), overlaid on the contour plot of $F_{x}^{RS}-F_{x}^{SCAN}$. There are considerable (s, q) points appear between $q_{0}(s)$ and $q_{H}(s)$ lines, and some points with $s>1.5$ are even above the $q_{0}(s)$ line, as shown in the inset plot. Given the quick decay of $F_{x}^{RS}$ vs. $q-q_{0}(s)$ shown in Figure 3 (b) and the exact match to the exchange energy of H, the above mentioned points have smaller $F_{x}$ values with RS than with SCAN, resulting in a less negative energy that matches the exact HF energy of $H_{2}^{+}$ at equilibrium as shown in Figure 1. As the bond length increases, for example at R = 2.116 Å, more (s, q) points appears above $q_{0}(s)$, as shown in section VII of Supplementary Material, thereby improving the energy over SCAN.
However, after R = 4 Å, the improvement starts to diminish and gets slightly worse than SCAN as shown in Figure 1. This is because the (s, q) points for $R=4.76$ appearing above $q_{0}(s)$ have too large q (q > 6) as shown in Figure 4(b), associated with relatively small s and very low electron densities (see section II of Supplementary Materials), and become energetically irrelevant. The energetically relevant electron density concentrates around the atomic centers with (s, q) below $q_{H}(s)$. This results in more negative exchange energy for RS than for SCAN because $F_{x}^{RS}$ decays slower than $F_{x}^{SCAN}$ with $s$ below $q_{H}(s)$. This is also dictated by the exchange hole of stretched bonds being too shallow, as shown in Figure 2.
In summary, we have developed the non-empirical $\nabla^{2}n$-dependent RS metaGGA that effectively reduces self-interaction errors in one-electron systems, while satisfying all exact constraints for one-electron systems and being guided by appropriate norms. The binding energy of $H_{2}^{+}$ predicted by RS matches the exact HF one beyond the equilibrium point, significantly improving over SCAN across a wide range of bond lengths. The inclusion of $\nabla^{2}n$ is shown to be crucial for reducing SIE of $H_{2}^{+}$ at equilibrium and stretched bonds. Our work therefore paves the way for further reducing SIE within the semilocal framework at the metaGGA level while maximally satisfying suitable exact constraints with guidance from appropriate norms.
Acknowledgments: A.R. and J.S. acknowledge the support of the U.S. Office of Naval Research (ONR) Grant No. N00014-22-1-2673. L.H. and J.V.B. were supported by the National Science Foundation (NSF) under Grant No. DMR-2042618, who carried out the calculations related to the exchange holes. The work at Los Alamos National Laboratory was carried out under the auspices of the U.S. Department of Energy (DOE) National Nuclear Security Administration under Contract No. 89233218CNA000001. It was supported by the LANL LDRD Program, and in part by the Center for Integrated Nanotechnologies, a DOE BES user facility, in partnership with the LANL Institutional Computing Program for computational resources. A.M.T. is grateful for support from the European Research Council under H2020/ERC Consolidator Grant “topDFT” (Grant No. 772259).